# Do LVLMs Understand Charts? Analyzing and Correcting Factual Errors in Chart Captioning

### v1

@@ Abstract
Recent advancements in large vision-language models (LVLMs) have led to significant progress in generating natural language descriptions for visual content and thus enhancing various applications. One issue with these powerful models is that they sometimes produce texts that are factually inconsistent with the visual input. While there has been some effort to mitigate such inconsistencies in natural image captioning, the factuality of generated captions for structured document images, such as charts, has not received as much scrutiny, posing a potential threat to information reliability in critical applications. This work delves into the factuality aspect by introducing a comprehensive typology of factual errors in generated chart captions. A large-scale human annotation effort provides insight into the error patterns and frequencies in captions crafted by various chart captioning models, ultimately forming the foundation of a novel dataset, Chocolate. Our analysis reveals that even state-of-the-art models, including GPT-4V, frequently produce captions laced with factual inaccuracies. In response to this challenge, we establish the new task of Chart Caption Factual Error Correction and introduce ChartVE, a model for visual entailment that outperforms proprietary and open-source LVLMs in evaluating factual consistency. Furthermore, we propose C2TFec, an interpretable two-stage framework that excels at correcting factual errors. This work inaugurates a new domain in factual error correction for chart captions, presenting a novel evaluation mechanism, and demonstrating an effective approach to ensuring the factuality of generated chart captions. [^1] The code and data will be made publicly available: https://github.com/khuangaf/CHOCOLATE.
@@
[…]
@@ 1 Introduction
Large vision-language models (LVLMs) have recently shown impressive capabilities in generating natural language descriptions of visual content like images, videos and charts Li et al. (2020, 2022); Wang et al. (2022); OpenAI (2023b); Google (2023a); Liu et al. (2023c); Wang et al. (2023). Chart captioning is particularly important for data analysts, business analysts, and journalists who rely on accurate chart interpretations for decision-making and reporting. However, no prior work have studied the factuality of the generated captions. Factuality, also known as the faithfulness or factual consistency between inputs and outputs, of generated captions is of critical importance when leveraging LVLMs for both practical applications and research advancements. In the domain of information dissemination, for instance, such models can be deployed to automatically generate descriptions for charts in news articles Liu et al. (2021), educational materials Fu et al. (2022), and on social media platforms Monteiro et al. (2017). Ensuring that these captions are factual directly impacts the credibility and reliability of the content provider.
@@
To better understand the factual errors produced by chart captioning models, we first introduce a typology of factual errors particularly relevant in the chart domain. Using this scheme, we conduct a large-scale human annotation study to analyze the distributions and frequencies of various error types, such as Value Error and Label Error, across captions produced by various chart captioning models, from task-specific fine-tuned models to LVLMs (see Table 1). The annotated samples are then categorized into three splits, Lvlm (Large-vision Language Models), Llm (Large Language Models), and Ft (Fine-tuned), based on the architecture and the scale of the underlying models that generate the captions, and form a dataset which we named Chocolate. With this dataset collected, we aim to answer three main research questions. First, are state-of-the-art chart captioning models able to produce factual captions? We find the answer is no (Section 2). Out of the 1,187 captions we analyzed, only 213 of them are factual (see Table 2). Even state-of-the-art LVLM, GPT-4V OpenAI (2023b), produces a great portion of errors in its generated captions (see Figure 1). In addition, we show that different chart captioning models exhibit substantially different factual error distributions in generated captions across different chart domains.
@@
The prevalence of factual inconsistencies observed in the generated captions by various models underscores the urgent need to address the factual accuracy of such technologies. Consequently, we introduce a task named Chart Caption Factual Error Correction (Section 3), which presents a novel challenge of rectifying factual inaccuracies in chart captions generated by LVLMs. A pertinent question that arises from this task is: how can we automatically evaluate the factual consistency between charts and captions/corrections? To tackle this question, we conceptualize the evaluation procedure as a task of visual entailment. We present the first model for chart visual entailment, named ChartVE. This model is trained by repurposing existing datasets from chart summarization and chart question answering. Our comprehensive correlation analysis shows that ChartVE performs competitively with proprietary LVLMs, GPT-4 and Bard Google (2023a), and outperforms an open-source LVLM, LLaVA-1.5 Liu et al. (2023c), despite being considerably smaller—100 times less in size. The findings emphasize the efficiency and efficacy of ChartVEn evaluating factual accuracy for the task of Chart Caption Factual Error Correction.
@@
Figure 1: Error distribution for different models on VisText and Pew. The error rates are computed per sentence. An error rate of 0.4 indicates that 40% of the sentences in the generated captions contain such an error. Note that a single caption may contain multiple error types; hence, the maximum value for a stacked bar is greater than 1.0. We show that even the most advanced LVLM, GPT-4, generates captions with a high rate of factual error.
@@
Now that we have set up the task, we turn to the challenge of how to effectively correct factual errors in chart captions? Here, we propose C2TFec (Section 4), an interpretable two-step framework that breaks down visual reasoning into image-to-structure rendering and text-based reasoning. C2TFec first transforms the input chart into a structured data table representation. Grounded in this extracted tabular data, the second component then identifies and fixes any factual inconsistencies in the initial caption through a transparent reasoning process. Our experiments demonstrate that this explicit decomposition enables more reliable factuality corrections compared to end-to-end approaches. The intermediate symbolic data acts as an effective bridge between charts and captions. By exploiting the complementary strengths of a specialized Chart-To-Table model and GPT-4, C2TFec significantly outperforms competitive baselines including GPT-4V, as validated through both automatic evaluation and human assessment.
@@
  • We present the first empirical analysis of factual errors in captions produced by models of various scales and architectures.
  • We introduce the Chart Caption Factual Error Correction task that challenges models to correct factual errors in generated chart captions and construct the Chocolate dataset for this task by defining a novel typology of factual errors in chart captions.
  • We present ChartVE, a reference-free evaluation metric based on visual entailment that correlates better with human judges compared to LVLM-based approaches.
  • We propose C2TFec, an interpretable two-stage error correction framework that demonstrates better performance compared to all existing LVLMs.
2 Dataset Creation
To understand the capabilities of existing models in summarizing key information from charts, we conduct a large-scale analysis on six most advanced chart captioning models on the VisText Tang et al. (2023a) and Pew Kantharaj et al. (2022) datasets. To facilitate this process, we introduce an error typology, as illustrated in Section 2.1. Upon gathering human annotations, we discuss the quality of the collected data (Section 2.3) and present a detailed analysis of different captioning models (Section 2.2).
@@ 2.1 Error Typology
Value Error
Asians have a turnout rate of 20.4% in 1990.
Label Error
Asians have the highest turnout rates across the years.
Trend Error
From 1986-2014, the turnout rates are increasing overall.
Magnitude Error
From 1986-2014, the turnout rates are sharply decreasing overall.
Out-of-context Error
Vietnamese have the highest turnout rates among Asians.
Nonsense Error
From 1986-2014, #?sep #sep #sep #sep.
Grammatical Error
The turnout rates are decrease overall.
@@
The overall direction of change over time or comparison between groups is incorrectly described in the caption, such as stating an increasing trend when the values in the chart are decreasing.
@@
The degree or amount of difference described for a trend or comparison is factually inconsistent with the chart, such as stating a small/sharp increase/decrease when the change is more gradual based on the chart’s data.
@@
Concepts, variables, or any information introduced in the caption that do not exist at all in the content of the chart. The caption contains factual statements not grounded in the actual chart contents.
@@
There are grammatical mistakes in the structure or syntax of the caption. [^2] Note that we do not consider grammatical errors as factual inconsistency. We analyze grammatical errors to understand how fluent the captioning models are.
@@ 2.2 Captioning Model Analysis
We consider the various types of models. First, ChartT5 Zhou et al. (2023), MatCha Liu et al. (2023b), and UniChart Masry et al. (2023) are the most advanced task-specific models fine-tuned with in-domain data from the VisText and Pew datasets. Second, DePlot + GPT-4 Liu et al. (2023a); OpenAI (2023a) is a LLM-based pipeline approach. Finally, GPT-4V OpenAI (2023b) and Bard Google (2023a) [^3] We tested Bard before the release of Gemini Google (2023b). are the strongest LVLMs. For each model and dataset, we randomly sample 100 chart figures and generate the corresponding captions. Invalid output sequences, such as empty strings, are filtered out.
@@
We compute the percentage of sentences with factual errors for different models and datasets, with a breakdown of the proportion of different error types. Here, we focus on the error distribution on the sentence level instead of the caption level since different models generate captions of different lengths. A sentence-level evaluation can help mitigate this discrepancy and facilitate a fairer comparison.
@@
From Figure 1, we made the following observations. First, SOTA chart captioning models often fail to produce factual captions. Additionally, as shown in Table 2, we calculated the percentage of non-factual captions, revealing that 82.05% of captions contain at least one factual error. More importantly, even advanced models like GPT-4V and Bard, which have demonstrated proficiency in a variety of vision-language tasks, produce factually incorrect captions 81.27% of the time, as recorded in Table 5. These findings highlight the inherent difficulties of chart understanding tasks and the limitations of SOTA vision-language models.
@@
Second, task-specific chart captioning models and LVLMs show opposite trends on the two datasets. Task-specific fine-tuned models, including ChartT5, MatCha, and UniChart, produce fewer errors on the VisText Tang et al. (2023a) dataset. Conversely, LVLMs, including GPT-4V and Bard, generate significantly fewer errors on the Pew dataset. The key distinctions on these datasets are two: (1) the vast majority of the charts in Pew datasets have corresponding numbers labeled on or next to each data point and (2) charts in VisText are associated with fewer complex structures compared to those in the Pew dataset. We hypothesize that LVLMs may be better at capturing the numbers labeled next to the data points in charts, whereas task-specific models are more capable of reading the exact value of the data points by aligning them to axes. We show an example to validate this hypothesis in Figure 6.
@@
Third, LVLMs cannot consistently outperform task-specific fine-tuned models. Although LVLMs are trained on data with a much larger scale using many more parameters, task-specific models may perform better than LVLMs when suitable pre-training objectives and model architectures are adopted. For example, on the VisText dataset, we see that UniChart achieves a much better performance than Bard and is comparable to GPT-4V in terms of producing more factual captions. This can be attributed to UniChart’s various pre-training objectives tailored for chart understanding, enabling it to better identify the relationship between data points within the charts from VisText.
@@
The dataset resulting from the analysis is named Chocolate (Captions Have Often Chosen Obfuscating Lies About The Evidence), where each instance consists of a chart, a generated chart caption, and error types labeled by human annotators. Drawing insights from Tang et al. (2023b) that factual errors produced by different kinds of models may be easier or more difficult to identify, we split Chocolate into three categories. First, the Lvlm split corresponds to captions produced by GPT-4V and Bard. Second, the Llm split contains captions generated by DePlot + GPT-4. Finally, the Ft split is associated with captions from ChartT5, UniChart, and MatCha. The details of each split are shown in Appendix C.
@@ 3.2 Reference-free Evaluation With Chart Visual Entailment
There was no established metric for evaluating the factual consistency between a chart and the corresponding chart caption. In addition, since our dataset does not contain annotated reference corrected captions [^4] We did not collect reference corrected captions due to the challenges of curating high-quality references through crowd-sourcing., text-based metrics such as SARI Xu et al. (2016) cannot be adopted. As a solution, we propose ChartVE, a reference-free evaluation metric based on chart visual entailment, as detailed in the following paragraphs.
@@
We formulate the task of chart caption factual inconsistency detection as a chart visual entailment task. Given a chart caption sentence $c$ and a chart $\mathcal{E}$, the task is to predict whether the relationship from $\mathcal{E}$ to $c$ as Entailment (i.e. factually consistent) or NotEntailment (factually inconsistent). The main challenge of learning a visual entailment model for this task is the lack of data. To overcome this challenge, we repurpose data from relevant tasks, such as chart QA and chart captioning, as positive samples. Then, we propose a table-guided negative data generation to produce negative samples. In the following paragraphs, we detail how each step is achieved.
@@
We consider datasets from two tasks that are closely related to the chart visual entailment task: chart question answering and chart captioning. The chart question answering task requires models to respond to questions regarding a specific chart, with each instance comprising a chart accompanied by a related question and answer pair in natural language. We utilize two datasets from chart question answering: ChartQA Masry et al. (2022) and PlotQA Methani et al. (2020a). To adapt these datasets for ChartVE, we apply a QA2Claim model Huang et al. (2023b) to transform the question-answer pairs into declarative statements. Paired with the original charts, these statements form positive instances (Entailment).
Meanwhile, chart captioning datasets include charts accompanied by descriptive captions, which typically consist of multiple sentences. Repurposing captions from VisText Tang et al. (2023a) and Chart-To-Text Kantharaj et al. (2022), we segment them into individual sentences. Each sentence is then paired with the relevant chart to create a positive instance. These adaptations allow us to repurpose existing datasets to create positive training examples for ChartVE.
@@
Creating negative training samples is a vital step for balancing our dataset, which we achieve by perturbing the positive instances grounded in the underlying data tables of the charts. For any given chart $\mathcal{E}_{i}$, we utilize its source data table $\mathcal{A}_{{\mathcal{E}_{i}}}$ as the basis for generating factual inconsistencies. Our approach involves scanning through each value within $\mathcal{A}_{{\mathcal{E}_{i}}}$ and identifying if it appears as a substring within the corresponding positive caption sentence $c^{+}_{i}$. When a match is found, we introduce a factual error by substituting that substring with a different value from the same column in $\mathcal{A}_{{\mathcal{E}_{i}}}$, yielding a value or label-error infused negative sentence $c^{-}_{i}$. This approach ensures $c^{-}_{i}$ is relevant but factually inconsistent to the corresponding chart $\mathcal{E}_{i}$.
To incorporate trend errors, we rely on a predefined lexicon of trend-specific terminology, such as words denoting “increase” or “decrease.” We then detect these terms within $c^{+}_{i}$ and systematically replace them with their antonyms to generate a trend-dissenting statement.
Additionally, we synthesize out-of-context errors by pairing chart $\mathcal{E}_{i}$ with an unrelated caption sentence $c^{+}_{j}$ from the pool, ensuring that $c^{+}_{j}$ corresponds to a different chart, where $i\neq j$. This form of error generation effectively emulates captions that introduce concepts or data not represented in the target chart, broadening the set of negative samples for our training corpus.
@@
Chocolate Model Lvlm Llm Ft LLaVA-1.5-13B 0.002 0.057 0.214 Bard -0.014 0.105 0.291 GPT-4V 0.157 0.205 0.215 DePlot + GPT-4 0.129 0.117 0.109 ChartVE (Ours) 0.178 0.091 0.215
@@
To evaluate the effectiveness of different methods in assessing the factuality of generated captions on the Chocolate dataset, we employ Kendall’s Tau Kendall (1938) to compute the correlation between these methods and human judgments. Given the absence of prior work specifically focused on building factual inconsistency detection models for chart captions, we compare our ChartVE metric with methods that are capable of performing zero-shot reasoning including DePlot + GPT-4 Liu et al. (2023a); OpenAI (2023a), Bard Google (2023a), GPT-4V OpenAI (2023b), and the leading open-source LVLM, LLaVA-1.5-13B Liu et al. (2023c). The prompts for using these models as evaluation metrics can be found in Appendix E.
The meta-evaluation results are summarized in Table 3. We observed that, overall, metrics exhibit the strongest correlation with human judgment on the Ft split and the weakest on the Lvlm split. This pattern aligns with expectations: the Ft captions are littered with more obvious mistakes, including a substantial number of out-of-context and nonsense errors, as detailed in Figure 1. Conversely, errors stemming from LVLMs often demand intricate inferences regarding the data points’ positions relative to the axes, thus presenting a more substantial challenge for detection. Importantly, our ChartVE metric demonstrates the highest correlation on the most demanding Lvlm split. However, ChartVE underperforms on the Llm split, potentially due to shifts in token distribution, as DePlot + GPT-4 occasionally employs table-centric terminology (e.g., “columns” and “entries”) absent from ChartVE’s training data. Despite this, ChartVE compares favorably to proprietary LVLMs and, notably, outstrips the performance of the SOTA open-source LVLM, LLaVA-1.5-13B, despite ChartVE being 100 times smaller in scale.
On the Llm and Ft splits, the leading metrics are Bard and GPT-4V, respectively. Yet, what stands out is that Bard presents a negative correlation on the Lvlm split, signaling that LVLMs may falter significantly in evaluating the factuality of chart captions under certain conditions. As a result, we advocate for leveraging the best-performing metric for each split for evaluation.

### v2

@@ Abstract
Advances in large vision-language models (LVLMs) have led to significant progress in generating natural language descriptions for visual contents. These powerful models are known for producing texts that are factually inconsistent with the visual input. While some efforts mitigate such inconsistencies in natural image captioning, the factuality of generated captions for structured visuals, such as charts, has not received as much scrutiny. This work introduces a comprehensive typology of factual errors in generated chart captions. A large-scale human annotation effort provides insight into the error patterns in captions generated by various models, ultimately forming the foundation of a dataset, Chocolate. Our analysis reveals that even advanced models like GPT-4V frequently produce captions laced with factual inaccuracies. To combat this, we establish the task of Chart Caption Factual Error Correction and introduce ChartVE, a visual entailment model that outperforms current LVLMs in evaluating caption factuality. Furthermore, we propose C2TFec, an interpretable two-stage framework that excels at correcting factual errors. This work inaugurates a new domain in factual error correction for chart captions, presenting a novel evaluation metric, and demonstrating an effective approach to ensuring the factuality of generated chart captions. The code and data as well as the continuously updated benchmark can be found at: https://khuangaf.github.io/CHOCOLATE/.
@@
Kung-Hsiang Huang1,2∗    Mingyang Zhou3     Hou Pong Chan4∗     Yi R. Fung1     Zhenhailong Wang1     Lingyu Zhang3     Shih-Fu Chang3     Heng Ji1 1University of Illinois Urbana-Champaign     2Salesforce AI Research 3Columbia University     4DAMO Academy, Alibaba Group 1{yifung2, wangz3, hengji}@illinois.edu   2kh.huang@salesforce.com 3{mz2974, lz2814, sc250}@columbia.edu   4houpong.chan@alibaba-inc.com
**footnotetext: Work was done while Kung-Hsiang was at UIUC and Hou Pong was at the University of Macau.
@@ 1 Introduction
Large vision-language models (LVLMs) have recently shown impressive capabilities in generating natural language descriptions of visual content like images, videos and charts OpenAI (2023b); Google (2023a); Liu et al. (2023c); Wang et al. (2023). Chart captioning is particularly important for data analysts, business analysts, and journalists who rely on accurate chart interpretations for decision-making and reporting. However, no prior work has studied the factuality [^1] Factuality is also known as the faithfulness or factual consistency between inputs and outputs of the generated captions. Given that factuality is vital for credibility in applications of chart captioning in news articles Liu et al. (2021), educational resources Fu et al. (2022), and social media Monteiro et al. (2017), examining the truthfulness of generated captions is a critical concern.
@@
To understand the factual errors in chart captioning models, we introduce a typology of factual errors for the chart domain. Using this scheme, we conduct a large-scale human annotation study to analyze the distributions of various error types, such as Value Error and Label Error, in captions from various models, from task-specific fine-tuned models to LVLMs (see Table 1). The annotated samples are then categorized into three splits, Lvlm (Large-vision Language Models), Llm (Large Language Models), and Ft (Fine-tuned Vision-language Models), based on the architecture and the scale of the underlying models, and form a dataset which we named Chocolate. With this dataset collected, we aim to answer three main research questions. First, are state-of-the-art chart captioning models able to produce factual captions? We find the answer is no (Section 2). Specifically, 82.06% of the generated captions are non-factual (see Table 2). Even state-of-the-art LVLMs like GPT-4V OpenAI (2023b) produce a great portion of errors in its generated captions (see Figure 1).
@@
The prevalence of factual inconsistencies observed in the generated captions by various models underscores the urgent need to mitigate the factual errors of such models. Hence, we introduce a new task, Chart Caption Factual Error Correction (Section 3), which presents a novel challenge of rectifying factual inaccuracies in chart captions generated by LVLMs. A pertinent question that arises from this task is: how to automatically evaluate the factual consistency between charts and captions? To tackle this question, we present ChartVE, novel visual entailment approach to assess the factual consistency of chart captions. This model is trained by repurposing existing resources from chart summarization and chart question answering. Results show that ChartVE performs competitively with proprietary LVLMs and outperforms the most advanced open-source LVLM, despite being 64 times less in size.
@@
Figure 1: Error distribution for different models on VisText and Pew. The error rates are computed per sentence. An error rate of 0.4 indicates that 40% of the sentences in the generated captions contain such an error. Note that a single caption may contain multiple types of errors; hence, the maximum value for a stacked bar is greater than 1.0. We show that even the most advanced LVLM, GPT-4V, generates captions with a high rate of factual error.
@@
Now that we have set up the task, we turn to the challenge of how to effectively correct factual errors in chart captions? We propose C2TFec (Section 4), an interpretable two-step framework that decomposes visual reasoning into image-to-structure rendering and text-based reasoning. C2TFec first transforms the input chart into a structured data table representation. Grounded in this extracted tabular data, the second component then identifies and fixes any factual inconsistencies in the generated caption through an interpretable reasoning process. Our experiments demonstrate that this explicit decomposition enables more reliable factuality corrections compared to end-to-end approaches. The intermediate symbolic representation acts as an effective bridge between charts and captions, enabling C2TFec to significantly outperform competitive baselines including GPT-4V (Section 6).
@@
  • We present the first analysis of factual errors in captions produced by models of various scales using a novel error typology, which results in the Chocolate dataset.
  • We introduce the Chart Caption Factual Error Correction task that challenges models to correct factual errors in generated chart captions.
  • We present ChartVE, a reference-free evaluation metric based on visual entailment that correlates better with human judges than LVLMs.
  • We propose C2TFec, an interpretable two-stage error correction framework that performs better than all existing LVLMs.
2 Analyzing Factual Errors
To understand the capabilities of existing models in summarizing key information from charts, we conduct a large-scale analysis on six most advanced chart captioning models on the VisText Tang et al. (2023a) and Pew Kantharaj et al. (2022) datasets. To facilitate this process, we introduce an error typology, as illustrated in Section 2.1. Upon gathering human annotations, we present a detailed analysis of different captioning models (Section 2.2) and discuss the quality of the collected data (Section 2.3).
@@ 2.1 Error Typology
[Uncaptioned image] Value Error Asians have a turnout rate of 20.4% in 1990.
Label Error Asians have the highest turnout rates across the years.
Trend Error From 1986-2014, the turnout rates are increasing overall.
Magnitude Error From 1986-2014, the turnout rates are sharply decreasing overall.
Out-of-context Error Vietnamese have the highest turnout rates among Asians.
Nonsense Error From 1986-2014, #?sep #sep #sep #sep.
Grammatical Error The turnout rates are decrease overall.
@@
The overall direction of change over time or comparison between groups is incorrectly described in the caption, such as stating an increasing trend when it is actually decreasing.
@@
The degree or amount of difference described for a trend is unfaithful to the chart, such as stating an increase “sharp” when the chart shows it is actually “smooth”.
@@
Concepts, variables, or any information introduced in the caption that does not exist at all in the content of the chart. The caption contains factual statements not grounded in the actual chart contents.
@@
There are grammatical mistakes in the structure or syntax of the caption. [^2] Note that we do not consider grammatical errors as factual inconsistency. They are analyzed for assessing fluency.
@@ 2.2 Captioning Model Analysis
We consider various types of models. First, ChartT5 Zhou et al. (2023), MatCha Liu et al. (2023b), and UniChart Masry et al. (2023) are the most advanced task-specific models fine-tuned with in-domain data from the VisText and Pew datasets. Second, DePlot + GPT-4 Liu et al. (2023a); OpenAI (2023a) is a LLM-based pipeline approach. Finally, GPT-4V and Bard [^3] We tested Bard before Gemini’s release Google (2023b). are the strongest LVLMs. For each model and dataset, we randomly sample 100 chart figures and generate the corresponding captions. Invalid output sequences, such as empty strings, are filtered out.
@@
We compute the percentage of sentences with factual errors for different models and datasets, with a breakdown of different error types. Error rates are computed at the sentence level instead of the caption level since different models generate captions of different lengths. A sentence-level evaluation helps mitigate this discrepancy and facilitates a fairer comparison.
@@
From Figure 1, we made the following observations. First, SOTA chart captioning models often fail to produce factual captions. Additionally, as shown in Table 2, we calculated the percentage of non-factual captions, revealing that 82.06% of captions contain at least one factual error. More importantly, even models like GPT-4V and Bard, which have demonstrated proficiency in a variety of vision-language tasks, produce factually incorrect captions 81.27% of the time, as recorded in Table 10. These findings highlight the inherent difficulties of chart captioning tasks and the limitations of SOTA vision-language models.
@@
Second, task-specific chart captioning models and LVLMs show opposite trends on the two datasets. Task-specific models, including ChartT5, MatCha, and UniChart, produce fewer errors on the VisText dataset. Conversely, LVLMs, including GPT-4V and Bard, generate significantly fewer errors on the Pew dataset. The key distinctions on these datasets are two: (1) the prevalent labeled values on charts from Pew and (2) the simpler structures in charts from VisText. We hypothesize that LVLMs may be better at utilizing the labeled numbers, while task-specific effectively interpret values via axis alignment. We show an example to validate this hypothesis in Figure 6.
@@
Third, LVLMs cannot consistently outperform task-specific fine-tuned models. Despite their extensive training data and parameters, LVLMs may be surpassed by task-specific models with appropriate pre-training objectives and architectures. For example, on the VisText dataset, UniChart outperforms Bard and is comparable to GPT-4V in terms of producing more factual captions owing to UniChart’s various pre-training objectives for chart comprehension, enabling better interpretation of the relationship between data points within charts.
@@
The dataset resulting from the analysis is named Chocolate (Captions Have Often ChOsen Lies About The Evidence), where each instance consists of a chart, a generated chart caption, and error types labeled by human annotators. Drawing insights from Tang et al. (2023b) that factual errors produced by different kinds of models may be easier or more difficult to identify, we categorize Chocolate into three splits: the Lvlm split, with captions from GPT-4V and Bard; the Llm split, featuring DePlot + GPT-4 outputs; and the Ft split, for ChartT5, UniChart, and MatCha captions. Split details are in Appendix C.
@@ 3.2 Reference-free Evaluation With Chart Visual Entailment
There was no established metric for evaluating the factual consistency between a chart and the corresponding chart caption. In addition, since our dataset does not contain annotated reference captions [^4] Reference captions are not collected due to the challenges of curating high-quality references through crowd-sourcing., text-based metrics cannot be adopted. As a solution, we propose ChartVE, a reference-free evaluation metric based on chart visual entailment, as detailed in the following paragraphs.
@@
We formulate the inconsistency detection problem as a chart visual entailment task. Given a chart caption sentence $c$ and a chart $\mathcal{E}$, the task is to predict whether the relationship from $\mathcal{E}$ to $c$ as Entailment (factually consistent) or NotEntailment (factually inconsistent). The main challenge of learning a visual entailment model for this task is the lack of data. To overcome this challenge, we repurpose data from relevant tasks, such as chart QA, as positive samples. Then, we propose a table-guided negative data generation to produce negative samples.
@@
We consider datasets from two tasks that are closely related to the chart visual entailment task: chart question answering and chart captioning. We utilize two datasets from chart question answering: ChartQA Masry et al. (2022) and PlotQA Methani et al. (2020). Using a QA2Claim model Huang et al. (2023b), we transform the question-answer pairs into declarative statements and pair them with the original charts to form positive instances (Entailment). For chart captioning, captions from VisText Tang et al. (2023a) and Chart-to-Text Kantharaj et al. (2022) are segmented into individual sentences. Each sentence is paired with the relevant chart to create a positive instance. These methods allow us to repurpose existing resources for training ChartVE.
@@
Generating negative training samples is achieved by perturbing the positive instances grounded in the underlying data tables of the charts. For a chart $\mathcal{E}_{i}$ and its underlying data table $\mathcal{A}_{{\mathcal{E}_{i}}}$, we locate values in $\mathcal{A}_{{\mathcal{E}_{i}}}$ that matches a substring within the positive caption $c^{+}_{i}$. When a match is found, the substring in the caption is substituted with a different value from the same column in $\mathcal{A}_{{\mathcal{E}_{i}}}$, yielding a value or label-error infused negative sentence $c^{-}_{i}$, maintaining relevance while ensuring inconsistency with $\mathcal{E}_{i}$. For trend-related errors, we replace trend-terms found in $c^{+}_{i}$ with their opposites, drawing on a specific lexicon of terms like “increase” and “decrease,” thereby creating trend-contradictory statements. Furthermore, out-of-context errors are crafted by pairing $\mathcal{E}_{i}$ with a mismatched caption $c^{+}_{j}$ from another chart, where $i\neq j$. This simulates captions filled with unrelated data.
@@
Chocolate Model Lvlm Llm Ft SummaC -0.011 0.023 0.036 QAFactEval 0.064 0.045 0.054 LLaVA-1.5-13B 0.002 0.057 0.214 ChartLlama 0.010 0.057 0.141 ChartAssistant-S 0.015 0.057 0.036 Bard -0.014 0.105 0.291 GPT-4V 0.157 0.205 0.215 DePlot + GPT-4 0.129 0.117 0.109 ChartVE (Ours) 0.178 0.091 0.215
@@
To evaluate the effectiveness of different methods in assessing the factuality of generated captions on the Chocolate dataset, we employ Kendall’s Tau Kendall (1938) to compute the correlation between these methods and human judgments. Given the absence of prior work on factual inconsistency detection methods for chart captions, we compare our ChartVE with zero-shot capable methods, including DePlot + GPT-4, Bard, GPT-4V, and the leading open-source LVLMs, LLaVA-1.5-13B Liu et al. (2023c), ChartLlama Han et al. (2023), and ChartAssistant-S Meng et al. (2024). Text-based factuality metrics, SummaC Laban et al. (2022) and QAFactEval Fabbri et al. (2022b), which compute the factual consistency between the reference caption and the generated caption, are also included. The prompts for these models are detailed in Appendix E.
Meta-evaluation, summarized in Table 3, shows that, overall, metrics exhibit the strongest correlation with human judgment on the Ft split and the weakest on the Lvlm split. This pattern aligns with expectations: the Ft captions are littered with more obvious mistakes, such as out-of-context and nonsense errors, while errors stemming from LVLMs are harder to detect since they often demand intricate inferences regarding the data points’ positions relative to the axes, as detailed in Figure 1. Importantly, Our ChartVE excels on the challenging Lvlm split, but less so on the Llm split, likely due to shifts in token distribution, as DePlot + GPT-4 occasionally employs table-centric terminology (e.g., “columns” and “entries”) absent from ChartVE’s training data. Despite this, ChartVE compares favorably to proprietary LVLMs and outperforms open-source LVLMs, despite ChartVE being 64 times smaller in scale.
Bard and GPT-4V lead on the Llm and Ft splits, respectively. However, Bard shows a negative correlation on the Lvlm split, hinting at LVLMs’ limitations in assessing the factuality of chart captions. Thus, we advocate for using the best-performing metric for each split for evaluation.